Family incomes have a mean of $60 000 with a standard deviation of $20 000. The data are normally distributed. Determine the probability that a randomly selected family has a family income less than $36 000. Include evidence of your work by typing out your solution fully.


1
Expert's answer
2021-08-04T11:18:45-0400

Solution:

Given:

"\mu=60000"

"\sigma=20000"

"P(X<36000)=P(Z<\frac{36000-60000}{20000})"

"=P(Z<-1.2)=P(Z>1.2)=1-P(Z\le1.2)
\\=1-0.88493
\\=0.11507"